OpenStudy (anonymous):

Sarah is the president of the drama club and wants to publish a newspaper. Expenses are $.90 for printing and mailing each copy, plus $600 total for research and writing. The price of the newsletter is $1.50 per copy. How many copies of the newsletter must the club sell to break even? a. To break even, the drama club must sell 1000 copies. b. To break even, the drama club must sell 100 copies. c. To break even, the drama club must sell 250 copies. d. To break even, the drama club must sell 444 copies.

OpenStudy (anonymous):

Where is the number at that you need to break even with

OpenStudy (anonymous):

I need the that # in order that I can help you

OpenStudy (anonymous):

Let n represent the number of centerpieces you sell. Write an equation to determine the cost, C. Write another equation to determine the revenue, R. .. (thats all it gives me )

OpenStudy (anonymous):

@Phebe

OpenStudy (phebe):

im here

OpenStudy (phebe):

what you need

OpenStudy (anonymous):

It's A

OpenStudy (phebe):

yes i think so too @sahrya is right

OpenStudy (anonymous):

0.90*1000+600=1500 = 1.50*1000=1500

OpenStudy (anonymous):

Get it?

OpenStudy (anonymous):

thanks !! :) both of you just saved my life haha

OpenStudy (phebe):

lol XD welcome

OpenStudy (anonymous):

You're welcome :)